Solving system using TI Nspire
2a-b=1
5a-3b=0

Answer by acw1213(28)

Not sure what you mean by TI Nspire.
You can use elimination here.
Multiply the first equation by 3.
3(2a - b = 1)
Your system is now:
6a - 3b = 3
5a - 3b = 0
Look at your coefficients. Are there any opposite signs or same signs?
YES! The "b"s have the same coefficients.
This means to ADD the equations.
The "b"s cancel out now.
11a = 3
Divide by 11 on both sides
a = 3/11
Plug in 3/11 for "a" in either equation and solve for "b"
5(3/11) - 3b = 0
15/11 - 3b = 0
Subtract 15/11 on both sides
-3b = -15/11
Divide by -3 on both sides.
b = 15/33